OpenStudy (anonymous):

log3 x=4 that is little number 3 my answer is 81 what is yours?

OpenStudy (lone001):

do you mean log3*x=4

OpenStudy (anonymous):

\[\log_{3}x=4 \]

OpenStudy (lone001):

81 you are true

OpenStudy (lone001):

x=3^4=81

OpenStudy (anonymous):

awesome!! thanx i have about 3 more to get checked if you could follow me

OpenStudy (lone001):

yeah